OpenStudy (anonymous):

For the function f(x) = 2 – 3x, find f(0). A. –3 B. –1 C. 2 D. 5

OpenStudy (anonymous):

put x=0 in f(x) , see what u will get

OpenStudy (akashdeepdeb):

If you are given f(x) = 2 – 3x This means f(a) = 2 – 3a f(b) = 2 – 3b f(anything) = 2 – 3(anything)

OpenStudy (anonymous):

f(0)

OpenStudy (anonymous):

b?

OpenStudy (akashdeepdeb):

Can you explain why you think it is B? :D

OpenStudy (anonymous):

the other guy said f(0)=2-3(0) =-1

OpenStudy (akashdeepdeb):

Are you sure 2 - 3(0) = -1 ? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it right

OpenStudy (akashdeepdeb):

3(0) means 3 times 0 which IS ZERO! :D Anything * (zero) = zero. :) So f(0) = 2 - 3(0) = 2 - 0 = 2. Understood? :) What matters is understanding why you get something right, rather than just getting it right. :)

OpenStudy (anonymous):

yes but i got it right right